would you show me how to work these problems please?
5/9-(-1/2)=
9 4/21 - 6 4/7=
2 1/3 x 3/4=
7/8x(-1/4)=
f+5/12=3/4

Answer by rfer(16322)

5/9-(-1/2)
5/9+1/2
10/18+9/18=19/18=1 1/8
------------------------------------------
9 4/21-6 4/7=
193/21-138/21=55/21=2 31/50
------------------------------
2 1/3x3/4=
7/3 x 3/4=21/12=1 9/12
--------------------------
you try the others